# A Holistic Optimization Framework for Energy Efficient UAV-assisted Fog Computing: Attitude Control, Trajectory Planning and Task Assignment

### v1

@@ I Introduction
Attitude control, trajectory planning, and task assignment policy in UAV-assisted fog computing are very important processes that significantly impact the overall efficiency of the system [9, 10]. Specifically, attitude control (by adjusting pitch, roll, and yaw) ensures a stable and precise orientation of the UAV during operation, which is a necessary condition to maintain a high quality of communication with IoT devices and other fog nodes [11]. In addition, a stable attitude facilitates the UAV to effectively perform computation and storage tasks as a fog node. On the other hand, trajectory planning of the UAV can reduce its power consumption by identifying the most efficient path to collect data and tasks based on the locations of IoT devices [12]. Finally, task assignment refers to the process of deciding whether a specific task should be handled locally by the IoT device, processed in the fog layer (including the UAV), or offloaded to the central cloud. Specific objectives, including minimizing latency, maximizing throughput, or optimizing energy efficiency, can be achieved by smartly assigning tasks to appropriate devices. Task assignment is often jointly optimized with resource allocation, where transmission resources such as power and bandwidth are distributed among different transmission pairs in the network, to facilitate the transmission process and improve the overall efficiency. The four processes are inherently linked in UAV-assisted fog computing. For example, a UAV’s attitude control would ensure that it maintains optimal orientations while following a planned trajectory or processing a task. Also, when deciding whether to offload a certain task to the fog or the cloud and how to allocate relevant transmission resources, the energy consumption and latency for a certain UAV to reach the proximity of the IoT initiating the task along a planned trajectory should also be taken into account.
@@
In this work, we revise and extend the model in [14] to better align with the practical scenarios of UAV-assisted fog computing. Particularly, we consider a three-dimensional network space (compared to a two-dimensional plane as in [14]) with varying terrain elevations. We also add the attitude control and optimal structure design modules for a quadrotor UAV, to ensure the stability of flights, prevent potential structural damage, and reduce the latency when the UAV ascends or descends by two newly proposed Active Disturbance Rejection Control (ADRC) strategies [15]. Compared to most existing similar studies, (e.g.,[16, 6, 7, 8]), which only considered the flight attitude at independent static points, we jointly consider planning the optimal trajectory and determining the attitudes, taking into account the extra consumption required for changing attitude along the trajectory. Our research is expected to provide new useful insights in UAV-assisted fog computing applications for improving the overall performance and efficiency.
@@ II Related Work
In trajectory planning and/or task assignment problems, one notable issue of applying ACS or PSO is that the algorithm may enter the deadlock state where one or more tasks wait endlessly for resources [35]. Zhou et al. [36] and Hou et al. [37] both proposed ACS-based algorithms with enhanced communication mechanisms to avoid deadlocks. However, while the proposed methods managed to reduce the likelihood of deadlocks, they did not eliminate the possibilities of such undesirable events. Another effort to overcome the deadlock problem is to establish a dead-corner table (or referred as a dead angle table[38, 39], which updates the path strength using a penalty function to avoid the deadlock. While this approach effectively overcomes the deadlock problem, it introduces additional overheads in maintaining the table and thus is less efficient.
@@
[38],[39] Trajectory Planning Enhanced ACS & Dead-corner table Avoid path deadlock
@@ III-A Network Structure and Components
The positions of $K$ MDs are randomly generated according to a Poisson Point Process (PPP), with the coordinate of the $j$th MD denoted by $\boldsymbol{M}_{j}=\left[x_{j},y_{j},z_{j}\right]$. For convenience, we discretize a continuous time horizon with length $T$ uniformly into $N$ timeslots, and thus each timeslot has a length of $L=\frac{T}{N}$. We assume that $N$ is sufficiently large, or equivalently, the timeslots are sufficiently short, such that the position of the UAV can be considered to be fixed within each timeslot. We use $[x(t),y(t),z(t)]$, where $t\in\{1,2,\cdots,N\}$, to denote the position of the UAV at the $t$th timeslot. The UAV is powered by a battery with a maximum capacity of $B_{c}$. The height and speed are restricted to not exceed $z^{\max}$ and $v^{\max}$, respectively, at all times.
@@
For the $i$th task from the $j$th MD, we denote the energy consumption and delay as $E_{ij}$ and $D_{ij}$, respectively. For each task, the quadrotor UAV needs to move sufficiently close to the MD that initiates the task through a planned trajectory, in order to receive and further process the task. As mentioned earlier, a task may be executed locally, in the fog layer by the UAV, or further offloaded to the data centers in the central cloud. The computing results of any offloaded task need to be transmitted back to the initiating MD. In this context, we define $x_{ij}(t)=(x_{ij}^{\text{MD}}(t),x_{ij}^{\text{UAV}}(t),x_{ij}^{\text{DC}}(t))$ as an array consisting of binary variables that indicate the specific execution location of the $i$th task from the $j$th MD during the $t$th timeslot. For ease of reference, we list the key notations related to the system model in Table II.
@@
$v(t)$ The UAV’s velocity vector at the $t$th timeslot
@@ III-C Rotor Propeller Design of Quadrotor UAV
where $P_{g}$ is the standard atmosphere, $M_{g}$ is the molecular weight of gas, $R_{g}$ is the gas constant, $T(t)$ is the outside ambient temperature and $z(t)$ is the height of flight at the location of the UAV. It is noteworthy that $M_{g}$ and $R_{g}$ are constants in the ideal gas state, and their values are given in Table V.
@@ III-E Energy Consumption
We consider that the total energy consumption includes the consumption for the transmission and computation during the task offloading process, as well as those for the movement of the UAV. Hereafter, we use the binary variables $x_{ij}^{\text{MD}}(t)$, $x_{ij}^{\text{UAV}}(t)$, and $x_{ij}^{\text{DC}}(t)$ to specify the assignment of the $i$th task from the $j$th MD at the $t$th timeslot. A value of $1$ indicates that the task is assigned to the corresponding location.
@@ III-E2 Energy consumption for transmission
$e_{ij}^{\text{TR}}(t)=$ $\left(x_{ij}^{\text{UAV}}(t)+x_{ij}^{\text{DC}}(t)\right)\left(\frac{p_{j}(t)s_{ij}(t)}{C_{j}(t)r_{(j,u)}(t)}+\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}\right)$ (11)
$+x_{ij}^{\text{DC}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}+\frac{p_{D}(t)s_{ij}(t)}{C_{j}(t)r_{(D,u)}(t)}\right),$
@@ III-E3 Energy consumption for computation
$e_{ij}^{\mathrm{COMP}}(t)=x_{ij}^{\text{UAV}}(t)\delta_{u}s_{ij}(t)c_{ij}\left(f(t)\right)^{2}+x_{ij}^{\text{MD}}(t)\delta_{j}s_{ij}(t)c_{ij}\left(f_{j}(t)\right)^{2},$ (12)
@@ III-F1 Transmission delay
$d_{ij}^{\text{TR}}(t)$ $=x_{ij}^{\text{DC}}(t)\left(\frac{s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}+\frac{s_{ij}(t)}{C_{j}(t)r_{(D,u)}(t)}\right)$ (14)
$+\left(x_{ij}^{\text{UAV}}(t)+x_{ij}^{\text{DC}}(t)\right)\left(\frac{s_{ij}(t)}{C_{j}(t)r_{(j,u)}(t)}+\frac{s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}\right).$
@@ III-F2 Computation delay
$d_{ij}^{\mathrm{COMP}}(t)=x_{ij}^{\text{UAV}}(t)\frac{s_{ij}(t)c_{ij}}{f(t)}+x_{ij}^{\text{MD}}(t)\frac{s_{ij}(t)c_{ij}}{f_{j}(t)}.$ (15)
@@ IV-B Joint optimization model
We consider that the operational efficiency cost of the network is determined by the weighted sum of the energy and delay components. For the $t$th timeslot, the objective function is defined for all tasks originating from the MDs, taking into account both total energy consumption defined in (13) and total time required to complete all tasks defined in (18). The joint optimization problem can then be defined as,
@@
$\begin{array}[]{ll}\underset{x_{ij}}{\operatorname{Minimize}}&\mathbb{S}(t)=\mathbb{D}(t)+\epsilon\cdot\mathbb{E}(t)\\ \text{S.t.}&\left(\complement_{1}\right)\hskip 5.69046pt0\leq v(t)\leq v^{\max}\\ &\left(\complement_{2}\right)\hskip 5.69046pt0\leq z(t)\leq z^{\max}\\ &\left(\complement_{3}\right)\hskip 5.69046pt\sum_{j=1}^{K}C_{j}(t)\leq N_{c}\\ &\left(\complement_{4}\right)\hskip 5.69046ptf^{\min}\leq f(t)\leq f^{\max}\\ &\left(\complement_{5}\right)\hskip 5.69046ptf_{j}^{\min}\leq f_{j}(t)\leq f_{j}^{\max}\\ &\left(\complement_{6}\right)\hskip 5.69046ptp^{\min}\leq p(t)\leq p^{\max}\\ &\left(\complement_{7}\right)\hskip 5.69046ptp_{j}^{\min}\leq p_{j}(t)\leq p_{j}^{\max}\\ &\left(\complement_{8}\right)\hskip 5.69046ptx_{ij}^{\text{MD}}(t)+x_{ij}^{\text{UAV}}(t)+x_{ij}^{\text{DC}}(t)=1\\ &\left(\complement_{9}\right)\hskip 5.69046ptx_{ij}^{\text{MD}}(t),x_{ij}^{\text{UAV}}(t),x_{ij}^{\text{DC}}(t)\in\{0,1\}\\ &\left(\complement_{10}\right)\sum_{t=0}^{T}\left(E^{\text{UAV}}(t)\right)\leq B_{c}\\ &i\in\{1,\ldots,N_{j}\},j\in\{1,\ldots,K\},\end{array}$ (21)
@@
The parameter $\epsilon$ is a weighting factor in the objective function. One can tune this weighting factor to adjust the relative importance of delay and energy consumption, flexibly catering to different magnitudes and ranges of the two measurements, as well as different requirements under various scenarios in UAV-assisted fog computing and reflecting the preference in terms of the trade-off between the two metrics. In terms of the constraints, $\complement_{1}$ to $\complement_{3}$ were described in detail previously when they first appeared. $\complement_{4}$ to $\complement_{7}$ ensure that the processing frequency and transmission power of UAV and MDs are always within the effective range. $\complement_{8}$ and $\complement_{9}$ guarantee that all tasks are at exactly one place during a timeslot. $\complement_{10}$ restricts that the total energy consumption of the UAV during the whole period of $T$ does not exceed the capacity of its battery, where $E^{\text{UAV}}(t)$ denotes the energy consumption of the UAV at the $t$th timeslot,
$E^{\text{UAV}}(t)$ $=E^{\text{MOV}}(t)+\sum_{j=1}^{K}\sum_{i=1}^{N_{j}}\left[x_{ij}^{\text{DC}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}\right)\right.$ (22)
$\left.+x_{ij}^{\text{UAV}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}+\delta_{u}s_{ij}(t)c_{ij}\left(f(t)\right)^{2}\right)\right].$
@@
Note that the processing frequencies $f(t)$, $f_{j}(t)$, and transmission powers $p(t)$, $p_{j}(t)$ are dependent on the decision variables $x_{ij}(t)$. Numerous scholarly works provide comprehensive analyses of these intricate interactions, e.g., [13].
@@ V-A PSO for Task Assignment and Resource Allocation
We first introduce an algorithm based on the idea of Particle Swarm Optimization (PSO), for assigning tasks and allocating resources given that the trajectory is already known. The steps of our proposed algorithm are demonstrated in Algorithm 1, which we refer the readers to the conference version of this paper [14] for more detailed explanations. We categorize the job assignment decision, transmission power, and processing frequency as a collective entity $s_{h}=(x_{ij},p_{j}(t),f_{j}(t))$ within the particle swarm. Here, $h\in\{1,...,H\}$. Next, the determination of the particle group’s position is carried out by uniformly sampling a set of particle groups denoted by $H\in\mathbb{N_{+}}$. In order to search for the optimal solution, the velocity of the particle groups is initialized and subsequently updated using the method in [41]. The algorithm terminates when the discrepancy between the outcomes of $20$ successive rounds is less than a specified threshold $\eth$, and then outputs the minimal value $\mathbb{S}^{*}$, the optimal decision for job assignment $x_{ij}^{*}$, the transmission power $p_{j}^{*}(t)$, and the processing frequency $f_{j}^{*}(t)$.
@@
17:     $(x_{ij}^{*}$; $p_{j}^{*}(t)$; $f_{j}^{*}(t))$ $\longleftarrow gBest$
18:Output $\mathbb{S}^{*}(t)$; $(x_{ij}^{*}$; $p_{j}^{*}(t)$; $f_{j}^{*}(t))$

### v2

@@ I Introduction
Attitude control, trajectory planning, and task assignment policy in UAV-assisted fog computing are very important processes that significantly impact the overall efficiency of the system [9, 10, 11]. Specifically, attitude control (by adjusting pitch, roll, and yaw) ensures a stable and precise orientation of the UAV during operation, which is a necessary condition to maintain a high quality of communication with IoT devices and other fog nodes [12]. In addition, a stable attitude facilitates the UAV to effectively perform computation and storage tasks as a fog node. On the other hand, trajectory planning of the UAV can reduce its power consumption by identifying the most efficient path to collect data and tasks based on the locations of IoT devices [13]. Finally, task assignment refers to the process of deciding whether a specific task should be handled locally by the IoT device, processed in the fog layer (including the UAV), or offloaded to the central cloud. Specific objectives, including minimizing latency, maximizing throughput, or optimizing energy efficiency, can be achieved by smartly assigning tasks to appropriate devices. Task assignment is often jointly optimized with resource allocation, where transmission resources such as power and bandwidth are distributed among different transmission pairs in the network, to facilitate the transmission process and improve the overall efficiency. The four processes are inherently linked in UAV-assisted fog computing. For example, a UAV’s attitude control would ensure that it maintains optimal orientations while following a planned trajectory or processing a task. Also, when deciding whether to offload a certain task to the fog or the cloud and how to allocate relevant transmission resources, the energy consumption and latency for a certain UAV to reach the proximity of the IoT initiating the task along a planned trajectory should also be taken into account.
@@
In this work, we revise and extend the model in [15] to better align with the practical scenarios of UAV-assisted fog computing. Particularly, we consider a three-dimensional network space (compared to a two-dimensional plane as in [15]) with varying terrain elevations. We also add the attitude control and optimal structure design modules for a quadrotor UAV, to ensure the stability of flights, prevent potential structural damage, and reduce the latency when the UAV ascends or descends by two newly proposed Active Disturbance Rejection Control (ADRC) strategies [16]. Compared to most existing similar studies, (e.g.,[17, 18, 6, 7, 8]), which only considered the trajectory planning and flight attitude at independent static points, we jointly consider planning the optimal trajectory and determining the attitudes, taking into account the extra consumption required for changing attitude along the trajectory. Our research is expected to provide new useful insights in UAV-assisted fog computing applications for improving the overall performance and efficiency.
@@ II Related Work
In trajectory planning and/or task assignment problems, one notable issue of applying ACS or PSO is that the algorithm may enter the deadlock state where one or more tasks wait endlessly for resources [37]. Zhou et al. [38] and Hou et al. [39] both proposed ACS-based algorithms with enhanced communication mechanisms to avoid deadlocks. While the proposed methods managed to reduce the likelihood of deadlocks, they did not eliminate the possibilities of such undesirable events. Another effort to overcome the deadlock problem is to establish a dead-corner table (or referred as a dead angle table[40], which updates the path strength using a penalty function to avoid the deadlock. While this approach effectively overcomes the deadlock problem, it introduces additional overheads in maintaining the table and thus is less efficient.
Another popular emerging approach to solve UAV-related planning and optimization problem is the deep reinforcement learning (DRL) (e.g., [41, 42]). However, we focus on classical heuristic algorithms (such as ACS and PSO), as they are known to be more scalable, more adaptive and less computationally intensive compared to DRL-based approaches [43]. These advantages make them particularly well-suited for the dynamic and large-scale nature of UAV-assisted fog computing scenarios.
@@
[9] Trajectory Planning & Task Assignment Convex Optimization Min. energy consumption and latency
@@
[44],[40],[45] Trajectory Planning Enhanced ACS & Dead-corner table Avoid path deadlock & Min. Task transfer latency
@@ III-A Network Structure and Components
The positions of $K$ MDs are randomly distributed according to a Poisson Point Process (PPP), with the coordinate of the $j$th MD denoted by $\boldsymbol{M}_{j}=\left[x_{j},y_{j},z_{j}\right]$. For convenience, we discretize a continuous time horizon with length $T$ uniformly into $N$ timeslots, and thus each timeslot has a length of $L=\frac{T}{N}$. We assume that $N$ is sufficiently large, or equivalently, the timeslots are sufficiently short, such that the position of the UAV can be considered to be fixed within each timeslot. We use $[x(t),y(t),z(t)]$, where $t\in\{1,2,\cdots,N\}$, to denote the position of the UAV at the $t$th timeslot. The UAV is powered by a battery with a maximum capacity of $B_{c}$. The height and speed are restricted to not exceed $z^{\max}$ and $v^{\max}$, respectively, at all times.
@@
For the $i$th task from the $j$th MD, we denote the energy consumption and delay as $E_{ij}$ and $D_{ij}$, respectively. For each task, the quadrotor UAV needs to move sufficiently close to the MD that initiates the task through a planned trajectory, in order to receive and further process the task. As mentioned earlier, a task may be executed locally, in the fog layer by the UAV, or further offloaded to the data centers in the central cloud. The computing results of any offloaded task need to be transmitted back to the initiating MD. In this context, we define $\mathbf{o}_{ij}(t)=(o_{ij}^{\text{MD}}(t),o_{ij}^{\text{UAV}}(t),o_{ij}^{\text{DC}}(t))$ as an array consisting of binary variables that indicate the specific execution location of the $i$th task from the $j$th MD during the $t$th timeslot. For ease of reference, we list the key notations related to the system model in Table II.
@@
$\mathbf{v}(t)$ The UAV’s velocity vector at the $t$th timeslot
@@ III-C Rotor Propeller Design of Quadrotor UAV
where $P_{g}$ is the standard atmosphere, $M_{g}$ is the molecular weight of gas, $R_{g}$ is the gas constant, $T(t)$ is the outside ambient temperature, $z(t)$ is the height of flight at the location of the UAV, and $M_{g}$ and $R_{g}$ are constants in the ideal gas state.
@@ III-E Energy Consumption
We consider that the total energy consumption includes the consumption for the transmission and computation during the task offloading process, as well as those for the movement of the UAV. Hereafter, we use the binary variables $o_{ij}^{\text{MD}}(t)$, $o_{ij}^{\text{UAV}}(t)$, and $o_{ij}^{\text{DC}}(t)$ to specify the assignment of the $i$th task from the $j$th MD at the $t$th timeslot. A value of $1$ indicates that the task is assigned to the corresponding location.
@@ III-E2 Energy consumption for transmission
$e_{ij}^{\text{TR}}(t)=$ $\left(o_{ij}^{\text{UAV}}(t)+o_{ij}^{\text{DC}}(t)\right)\left(\frac{p_{j}(t)s_{ij}(t)}{C_{j}(t)r_{(j,u)}(t)}+\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}\right)$ (11)
$+o_{ij}^{\text{DC}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}+\frac{p_{D}(t)s_{ij}(t)}{C_{j}(t)r_{(D,u)}(t)}\right),$
@@ III-E3 Energy consumption for computation
$e_{ij}^{\mathrm{COMP}}(t)=o_{ij}^{\text{UAV}}(t)\delta_{u}s_{ij}(t)c_{ij}\left(f(t)\right)^{2}+o_{ij}^{\text{MD}}(t)\delta_{j}s_{ij}(t)c_{ij}\left(f_{j}(t)\right)^{2},$ (12)
@@ III-F1 Transmission delay
$d_{ij}^{\text{TR}}(t)$ $=o_{ij}^{\text{DC}}(t)\left(\frac{s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}+\frac{s_{ij}(t)}{C_{j}(t)r_{(D,u)}(t)}\right)$ (14)
$+\left(o_{ij}^{\text{UAV}}(t)+o_{ij}^{\text{DC}}(t)\right)\left(\frac{s_{ij}(t)}{C_{j}(t)r_{(j,u)}(t)}+\frac{s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}\right).$
@@ III-F2 Computation delay
$d_{ij}^{\mathrm{COMP}}(t)=o_{ij}^{\text{UAV}}(t)\frac{s_{ij}(t)c_{ij}}{f(t)}+o_{ij}^{\text{MD}}(t)\frac{s_{ij}(t)c_{ij}}{f_{j}(t)}.$ (15)
@@ IV-B Joint optimization model
We consider that the operational efficiency cost of the network is determined by the weighted sum of the energy and delay components. Therefore, our objective function consists of the energy consumption defined in (13) and the time required to complete all tasks defined in (18), during the entire period of $T$, as follows.
@@
$\begin{array}[]{ll}\underset{\{\mathbf{o}_{ij}(t),\mathbf{v}(t),\omega_{i}(t)|t\in\{0,\cdots,T\}\}}{\operatorname{Min}}&\mathbb{S}=\sum\limits_{t=0}^{T}[\mathbb{D}(t)+\epsilon\cdot\mathbb{E}(t)]\\ \text{s.t.}&\left(\complement_{1}\right)\hskip 5.69046pt0\leq|\mathbf{v}(t)|\leq v^{\max}\\ &\left(\complement_{2}\right)\hskip 5.69046pt0\leq z(t)\leq z^{\max}\\ &\left(\complement_{3}\right)\hskip 5.69046pt\sum_{j=1}^{K}C_{j}(t)\leq N_{c}\\ &\left(\complement_{4}\right)\hskip 5.69046ptf^{\min}\leq f(t)\leq f^{\max}\\ &\left(\complement_{5}\right)\hskip 5.69046ptf_{j}^{\min}\leq f_{j}(t)\leq f_{j}^{\max}\\ &\left(\complement_{6}\right)\hskip 5.69046ptp^{\min}\leq p(t)\leq p^{\max}\\ &\left(\complement_{7}\right)\hskip 5.69046ptp_{j}^{\min}\leq p_{j}(t)\leq p_{j}^{\max}\\ &\left(\complement_{8}\right)\hskip 5.69046pto_{ij}^{\text{MD}}(t)+o_{ij}^{\text{UAV}}(t)+o_{ij}^{\text{DC}}(t)=1\\ &\left(\complement_{9}\right)\hskip 5.69046pto_{ij}^{\text{MD}}(t),o_{ij}^{\text{UAV}}(t),o_{ij}^{\text{DC}}(t)\in\{0,1\}\\ &\left(\complement_{10}\right)\sum_{t=0}^{T}\left(E^{\text{UAV}}(t)\right)\leq B_{c}\\ &i\in\{1,\ldots,N_{j}\},j\in\{1,\ldots,K\}\end{array}$ (21)
@@
The decision variables include the angular velocities $\omega_{i}(t)$ related to attitude control, the velocity vector $\mathbf{v}(t)$ for trajectory planning, and task assignment decisions $\mathbf{o}_{ij}(t)$. The parameter $\epsilon$ is a weighting factor in the objective function, which flexibly caters to different magnitudes and ranges of the two measurements, as well as reflects the preference in terms of the trade-off between the two metrics in various fog computing scenarios. We will demonstrate in the result section that our proposed approach can achieve high performances for a wide range of $\epsilon$ values. In terms of the constraints, $\complement_{1}$ to $\complement_{3}$ were described in detail previously when they first appeared. $\complement_{4}$ to $\complement_{7}$ ensure that the processing frequency and transmission power of UAV and MDs are always within the effective range. $\complement_{8}$ and $\complement_{9}$ guarantee that every task is executed at exactly one place at any moment (no duplicate computing). $\complement_{10}$ restricts that the total energy consumption of the UAV during $T$ does not exceed its battery capacity, as $E^{\text{UAV}}(t)$ denotes the energy consumption of the UAV at the $t$th timeslot,
$E^{\text{UAV}}(t)$ $=E^{\text{MOV}}(t)+\sum_{j=1}^{K}\sum_{i=1}^{N_{j}}\left[o_{ij}^{\text{DC}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,D)}(t)}\right)\right.$ (22)
$\left.+o_{ij}^{\text{UAV}}(t)\left(\frac{p_{u}(t)s_{ij}(t)}{C_{j}(t)r_{(u,j)}(t)}+\delta_{u}s_{ij}(t)c_{ij}\left(f(t)\right)^{2}\right)\right].$
@@
Note that the processing frequencies $f(t)$, $f_{j}(t)$, and transmission powers $p(t)$, $p_{j}(t)$ are dependent on the decision variables $\mathbf{o}_{ij}(t)$. Numerous scholarly works provide comprehensive analyses of these intricate interactions, e.g., [14].
@@ V-A PSO for Task Assignment and Resource Allocation
We first introduce an algorithm based on the idea of Particle Swarm Optimization (PSO), for assigning tasks and allocating resources given that the trajectory is already known. The steps of our proposed algorithm are demonstrated in Algorithm 1, which we refer the readers to the conference version of this paper [15] for more detailed explanations. We categorize the job assignment decision, transmission power, and processing frequency as a collective entity $s_{h}=(\mathbf{o}_{ij},p_{j}(t),f_{j}(t))$ within the particle swarm. Here, $h\in\{1,...,H\}$. Next, the determination of the particle group’s position is carried out by uniformly sampling a set of particle groups denoted by $H\in\mathbb{N_{+}}$. In order to search for the optimal solution, the velocity of the particle groups is initialized and subsequently updated using the method in [47]. The algorithm terminates when the discrepancy between the outcomes of $20$ successive rounds is less than a specified threshold $\eth$, and then outputs the minimal value $\mathbb{S}^{*}$, the optimal decision for job assignment $\mathbf{o}_{ij}^{*}$, the transmission power $p_{j}^{*}(t)$, and the processing frequency $f_{j}^{*}(t)$.
@@
17:     $(\mathbf{o}_{ij}^{*}$; $p_{j}^{*}(t)$; $f_{j}^{*}(t))$ $\longleftarrow gBest$
18:Output $\mathbb{S}^{*}(t)$; $(\mathbf{o}_{ij}^{*}$; $p_{j}^{*}(t)$; $f_{j}^{*}(t))$